OpenStudy (jordanloveangel):

1) Solve. 8x = 4x – 32 A. –4 B. –8 <----- C. 4 D. 8 2)What is the value of b? 11+4b=8b-13 b=? 3)What is the value of w? 2(2w-6)=6(w-6)+14 w=? 4)Which equation would best help solve this problem? Five added to 2 times a number is equal to 7 less than 4 times the number. A. 2x – 5 = 4x + 7 B. 5x + 2 = 7x – 4 C. 2x + 5 = 4x – 7 <----- D. 4x +7 = 2x – 5 5)Eight subtracted from 4 times a number is equal to 4 less than 6 times the number. What is the number?

OpenStudy (jordanloveangel):

@itsbribro can u help me please

OpenStudy (itsbribro):

2. b=6

OpenStudy (itsbribro):

for number 1 is is negative 32 or minus 32

OpenStudy (itsbribro):

3. w=5

OpenStudy (jordanloveangel):

for 1 is it a,b,c or d?

OpenStudy (itsbribro):

1. B -8

OpenStudy (jordanloveangel):

oh kk what about 4 and 5?

OpenStudy (itsbribro):

they have diamonds

OpenStudy (jordanloveangel):

Which equation would best help solve this problem? Five added to 2 times a number is equal to 7 less than 4 times the number. A. 2x – 5 = 4x + 7 B. 5x + 2 = 7x – 4 C. 2x + 5 = 4x – 7 D. 4x +7 = 2x – 5

OpenStudy (jordanloveangel):

hold on one sec

OpenStudy (jordanloveangel):

Which equation would best help solve this problem? Five added to 2 times a number is equal to 7 less than 4 times the number. A. 2x - 5=4x +7 B. 5x + 2=7x -4 C. 2x + 5 =4x -7 D. 4x +7 =2x -5

OpenStudy (itsbribro):

c

OpenStudy (itsbribro):

i thought it was minus

OpenStudy (jordanloveangel):

5. Eight subtracted from 4 times a number is equal to 4 less than 6 times the number. What is the number?

OpenStudy (itsbribro):

i think your equation would be (4x)-8=(6x)-4 or 8-4*x=4-6*x what do you think i suck at writeint an equation from words

OpenStudy (jordanloveangel):

yeah its (4x)-8=(6x)-4

OpenStudy (itsbribro):

ok lol

OpenStudy (itsbribro):

then x would = -2

OpenStudy (jordanloveangel):

oh kk thanx

OpenStudy (itsbribro):

your welcome medals will be awesome if you can :) and thanks

OpenStudy (jordanloveangel):

i did